I am stuck on this word problem and wondering if you would be able to assist me.

J. Bond and E. Blofeld are driving speedboats in opposite directions after a failed attempt by the latter to blow up the world with a nuclear bomb. Mr. Bond’s boat is going 25 MPH faster than Mr. Blofeld’s. After 3 hours they are 500 miles apart. How fast was each of them going?
joe

Answer by rfer(16322)

(x+x+25)(3)=500
(2x+25)(3)=500
6x+75=500
6x=425
x=70.8 mph
x+25=95.8 mph